# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v3

@@ Conclusion.
Latent encodings resulting from the DDIM Inversion deviate from the Gaussian noise towards zero values. This can be attributed to the less diverse values for parts of the latents corresponding to the plain image surfaces. This observation reveals that latent encodings retain some characteristics of the original input samples and contain information about the source generation.
@@ 3.3 Origin of the divergence
Figure 4: Inversion approximation error is significantly higher during the first steps of the inversion process than for the next steps. Moreover, in the first steps, errors for plain areas contribute more significantly to the error than errors for non-plain pixels.
Pixel area Diffusion steps Model
ADM-64 DiT
Plain $1,2\dots 50$ $100\%$ $15.11$ $5.67$
Non-plain $12.43$ $3.16$
Plain $1,2\dots 5$ $10\%$ $9.23$ $3.41$
Non-plain $6.80$ $1.42$
Plain $6,7\dots 50$ $90\%$ $5.87$ $2.23$
Non-plain $5.63$ $1.74$
Table 6: Total per-pixel inversion error (normalized) over different timestep ranges ¯$\bar{\mathcal{E}}_{(t_{s},t_{e})}^{E}$ for plain and non-plain areas. Inversion approximation error is higher for pixels related to plain image areas, especially in the first 10% of inversion steps.
@@
We first recall that the DDIM Inversion error can be attributed to the approximation of the diffusion model’s output at step $t\in{1\dots T}$ with the output from step $t-1$ (see eq. 3). Hence, we can define the inversion approximation error for step $t$ as the difference between DM’s output for the target and previous timesteps ${t}$ and ${t-1}$ as:
$\xi(t)=\lvert\underbrace{\epsilon^{(t)}_{\theta}(x_{t},c)}_{\mathcal{E}_{t}^{I}}-\underbrace{\epsilon^{(t)}_{\theta}(x_{t-1},c)}_{\mathcal{E}_{t}^{S}}\rvert,$ (4)
@@
where $\mathcal{E}_{t}^{S}$ is the true model prediction at step $t$, and $\mathcal{E}_{t}^{I}$ is the inversion’s approximation using the previous step’s output. Based on the observations from the previous section, we propose to investigate how the inversion approximation error $\xi(t)$ differs for pixels associated with plain and non-plain image areas throughout the inversion process. To that end, we average the approximation errors for $4000$ images for each of the $T=50$ diffusion timesteps. To ascertain that we measure the error solely for the analyzed step $t$, we start the inversion procedure from the exact latent from step $(t-1)$ (obtained from the sampling path). We split the latent pixels into plain and non-plain areas according to the masks calculated for clean images. More details on the experimental setup can be found in appendix H.
@@
In fig. 4, we present the visualization of calculated differences for each diffusion step. There is a significant difference in the prediction errors for plain and non-plain areas, especially in the initial steps of the noising process. Additionally, in table 6, we show the actual values of the approximation error. By splitting the inversion process into two parts, we can observe that the first 10% of inversion steps contribute more to the overall error than the remaining 90%. Notably, the error in plain areas predominantly accumulates within the first 10% of the inversion steps.
@@
Those observations can be related to recent works [29, 32] that analyze why numerical solvers incur significant errors during the earliest diffusion steps (as $t\rightarrow 0$). Specifically, Lee et al. [29] trace the error to the $1/t$ curvature blow-up of the reverse-time ODE trajectory, whereas Lin et al. [32] attribute the predominance of early DDIM inversion approximations to a singularity arising from commonly used noise schedules. Our experiments extend those studies by showing that the approximation error can mainly be attributed to the plain regions in the original images.
@@ Conclusion.
Early approximations during the DDIM Inversion procedure lead to unequally distributed errors for pixels related to the plain and non-plain image areas, making it the origin of structural patterns and correlations exhibited by the latent encodings.
4 Consequences of the divergence and how to fix them?
After identifying the differences between the Gaussian noises and latents and highlighting the origin of this phenomenon, we finally pose the last question:
Research Question 4: What are the practical consequences of the divergence between noises and inverse DDIM latents, and how to fix them?
Building on top of our findings in section 3.3, which indicated that the initial inversion steps predominantly contribute to the divergence between latent variables and Gaussian noise, we propose to decorrelate the final DDIM latents by replacing the first inversion steps with random Gaussian noise $\epsilon\sim\mathcal{N}(0;\mathcal{I})$ as in the standard forward diffusion process (see appendix I for pseudocode):
$x_{t}=\begin{cases}\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,&\text{if }t\leq t^{\prime}\quad\text{(Forward Diffusion)}\\ \sqrt{\alpha_{t}}{x_{t-1}}+(\sqrt{1-\bar{\alpha}_{t}}-\sqrt{\alpha_{t}-\bar{\alpha}_{t}})\cdot\epsilon^{(t)}_{\theta}(x_{t-1},c),&\text{if }t>t^{\prime}\quad\text{(DDIM Inversion)}\end{cases}$ (5)
We first evaluate the effectiveness of this technique, with $N=10000$ images generated with $T=50$ diffusion steps using DiT and IF models. For such generations, we first noise them with a forward diffusion to the intermediate step $t^{\prime}$, followed by the DDIM inversion for $T-t^{\prime}$ steps, starting from the $x_{t^{\prime}}$ and ending up with a new approximation of the starting noise. In table 7, we show that by replacing just $4\%$ of the inversion steps, we can completely remove correlations in the latents, up to the level of random Gaussian noise. Moreover, in table 8, we measure the size of the different parts of images (plain vs non-plain) after saving them with the PNG lossless compression. As is visible, compression is the most effective in the parts related to plain images generated from the DDIM latents, which inclines low diversity of their values. At the same time, replacing only 4% of inversion steps significantly reduces this issue. In the following sections, we showcase several consequences of the divergence between noises and latents and how our simple solution reduces those issues.
@@ 4 Consequences of the divergence and how to fix them?
forward (%$\%T$) Corr. Absolute Error ×$\times 10^{-3}$ Corr. Absolute Error ×$\times 10^{-3}$
@@ 4.1 Interpolation quality
We start with the task of image interpolation, where the goal, for two given images, is to generate a sequence of semantically meaningful intermediary frames. Numerous methods [8, 50, 47, 63, 60, 1] leverage diffusion models to obtain realistic transitions. Song et al. [50] propose to use the latent variables resulting from the DDIM inversion technique and interpolate them with the spherical linear interpolation (SLERP [48]), that, for two objects $x$ and $y$, with an interpolation coefficient $\lambda\in[0;1]$, is defined as $z^{(\lambda)}=\frac{\sin{(1-\lambda)\theta}}{\sin{\theta}}x+\frac{\sin{\lambda\theta}}{\sin{\theta}}y$, where $\theta=\arccos\left((x\cdot y)/(\lVert x\rVert\lVert y\rVert)\right)$.
We compare the quality of interpolations in the noise and latent spaces. To this end, we sample $N=20$k noises, use DDIM with $T=50$ diffusion steps to generate images, and invert those images back into the latents. Next, we randomly assign pairs of either two noises or latents, which we interpolate with SLERP for $\lambda\in\{0,\frac{1}{6},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{5}{6},1\}$ and denoise. In fig. 5, we show that, by calculating FID-10k, intermediate noise generations preserve consistent quality along the entire interpolation path – meaning that all interpolated images fall into the real images manifold. In contrast, this property collapses for the setting with latents as inputs. In the following rows, we show that worse interpolation results for latents stem from the decline in the variety of produced generations, as indicated by a lower recall [27], especially in the middle of the interpolation path. Nevertheless, as presented in table 9, replacing only the first step of the DDIM inversion with the forward diffusion significantly reduces the issue, leading almost to the quality of the interpolations between ground-truth noises.
Figure 5: The quality of images generated from spherical linear interpolations of latents ${\mathbf{\hat{x}_{T}}}$ deteriorates along the interpolation path, as indicated by the FID peak in the middle. In contrast, the quality of generations from noise ${\mathbf{x_{T}}}$ interpolations remains stable. Recall metric reveals that the latent interpolation leads to the generation of much less diversity than noise.
@@ 4.2 Diversity and quality of image edition
Table 9: Starting inversion with forward diffusion improves quality (FID) of latent interpolations. Replacing first steps improves generations’ diversity (Recall).
Inversion steps DiT
replaced by SLERPλ=0.5
forward (%$\%T$) FID $\downarrow$ Precision $\uparrow$ Recall $\uparrow$
Noise ${\mathbf{x_{T}}}$ $16.61$ $0.799$ $0.473$
DDIM Latent ${\mathbf{\hat{x}_{T}}}$ $19.68$ $0.795$ $0.415$
$1$ ($2\%$) $16.89$ $0.816$ $0.443$
$1\dots 2$ ($4\%$) $16.74$ $0.817$ $0.443$
$1\dots 5$ ($10\%$) $17.05$ $0.812$ $0.444$
$1\dots 10$ ($20\%$) $17.02$ $0.809$ $0.454$
Text-to-image diffusion models with DDIM inversion are often used to modify images [14, 38, 12, 19]. However, knowing that DDIM latents are less diverse in plain areas, we hypothesize that using them as a starting point might reduce the diversity and quality of the edited samples. To evaluate this, we leverage DiT [42] and IF [52] as conditional DMs with $T=50$ diffusion steps. For each model, we construct two sets of $1280$ randomly selected prompts (ImageNet classes for DiT, COCO-30K [31] prompts for IF): (1) source prompts $P_{S}$, used during the generation and inversion, and (2) target prompts $P_{T}$, used for edition. Using source prompts $P_{S}$, we generate images $I_{S}$ from Gaussian noise ${\mathbf{x_{T}}}$ and invert them back into the latents ${\mathbf{\hat{x}_{T}}}$. Next, we regenerate images $\hat{I}_{T}$ from the latents, changing the conditioning to the target prompts $P_{T}$. We compare the edits with ground truth targets $I_{T}$ generated from the original noise ${\mathbf{x_{T}}}$ with $P_{T}$. In fig. 6, we present the drawback of leveraging latent variables as starting points for the denoising, where the $I_{S}$ images’ structures in the latents limit editing performance at those locations in $\hat{I}_{T}$ images.
Figure 6: Image edition by reversing the original image with a source prompt and reconstructing it with a target one. DDIM Inversion produces less diverse image changes when the manipulation is related to plain regions in source images, contrary to when using ground-truth Gaussian noise. Replacing the first steps of DDIM Inversion with forward diffusion leads to more editable latents.
In table 10, we quantitatively measure this effect. First, we calculate the diversity of target generations ($I_{T}$,$\hat{I}_{T}$) against source images ($I_{S}$). We use DreamSim distance [11], LPIPS [61], SSIM [57], and cosine similarity of DINO features [5] to measure the distance between two sets of generations. The experiment shows that edits resulting from latent encodings $\hat{I}_{T}$ are characterized by higher similarity (SSIM, DINO) and lower diversity (DreamSim, LPIPS) relative to starting images $I_{S}$ than the one resulting from noises $I_{T}$. At the same time, in the bottom rows of table 10,
@@
against $I_{S}$ LPIPS $\uparrow$ $0.59_{\pm 0.12}$ $0.57_{\pm 0.12}$ $0.38_{\pm 0.11}$ $0.33_{\pm 0.11}$
@@
alignment CLIP-T ($P_{T}$) $\uparrow$ $0.689$ $0.568$ $0.353$ $0.614$
@@
Table 10: Diversity of editions (generations from noise ${\mathbf{x_{T}}}$ and latents ${\mathbf{\hat{x}_{T}}}$, conditioned on target prompt) in relation to source images $I_{S}$ and their alignment with source and target prompts. The arrows ($\uparrow$/$\downarrow$) indicate greater generation diversity and higher text alignment to the target prompt.
@@
we show that the correlations occurring in the latent encodings induce lower performance in text-alignment to target prompts $P_{T}$, which we measure by calculating cosine similarity between text embeddings and resulting image embeddings, both obtained with the CLIP [44] encoder. For the DiT model, our experiment indicates that generations from latents are as similar to the target prompt from the regeneration as to the source one.
@@
Figure 7: Replacing first DDIM inversion steps with forward diffusion improves editions’ pixel diversity (a) and prompt-alignment (b). For IF and DiT models, diversity of generations can be improved by leveraging the forward diffusion process in the first inversion steps, and denoising resulting latents with a different prompt. Additionally, we observe a boost in generations’ alignment to the target prompts, as indicated by the larger CLIP-T value for $P_{T}$ and the smaller one for $P_{S}$. We present details of the experiment in appendix J.
@@
We further evaluate how replacing the first inversion steps with the forward diffusion affects the diversity and text alignment of generated outputs. As shown in fig. 7a, swapping the first $4\%$ of DDIM inversion steps with forward diffusion improves the diversity of images generated from latents almost to the level of the samples from original noise. At the same time, as presented in fig. 7b, replacing the first inversion steps leads to a significant decrease in generations’ alignment to the source prompt ($P_{S}$) and an increase in similarity to the target prompt ($P_{T}$). This can also be observed in visualisations (fig. 6).
Conversely, replacing 4% of model predictions with a forward diffusion makes the process faster by around $3.9\%$. Although at the same time, intuitively, replacing the predicted noise with a random one should degrade reconstruction quality, recent studies [6, 33, 30, 10] have shown that the backward diffusion final steps do not contribute additional generative information, functioning instead as a data-agnostic denoising process. Therefore, substituting noise only in the initial steps of the inversion (e.g., the first 4%) allows us to enhance the quality of the latent representations while maintaining a reconstruction error at the level of the standard DDIM inversion. This means that the error between the ground truth prediction and DDIM approximation (eq. 4) is comparable to the error introduced with random Gaussian noise of the adjusted magnitude. Additionally, as presented in appendix K, our approach enables stochastic image editing, producing diverse manipulations of the same image.
On top of the experiments presented in this section, in the appendices, we provide additional studies comparing noise-to-image and latent-to-image mappings (appendix L) and the relation between noise and samples during training (appendix M).

### v4

@@ Conclusion.
Latent encodings resulting from the DDIM Inversion deviate from the Gaussian noise towards zero values. This is especially true for parts of the latents corresponding to the plain image surfaces. This observation reveals that latent encodings retain some characteristics of the original input samples and contain information about the source generation.
@@ 3.3 Origin of the divergence
We first recall that the DDIM Inversion error can be attributed to the approximation of the diffusion model’s output at step $t\in{1\dots T}$ with the output from step t1t-1 (see Eq. 3). Hence, we can define the inversion approximation error for step $t$ as the difference between DM’s output for the target and previous timesteps t{t} and t1{t-1} as:
$\xi(t)=\lvert\underbrace{\epsilon^{(t)}_{\theta}(x_{t-1},c)}_{\mathcal{E}_{t}^{I}}-\underbrace{\epsilon^{(t)}_{\theta}(x_{t},c)}_{\mathcal{E}_{t}^{S}}\rvert$, (4)
@@
where $\mathcal{E}_{t}^{S}$ is the true model prediction at step $t$, and $\mathcal{E}_{t}^{I}$ is the inversion’s approximation using the previous step’s output. Based on the observations from the previous section, we propose to investigate how the inversion approximation error $\xi(t)$ differs for pixels associated with plain and non-plain image areas throughout the inversion process. To that end, we average the approximation errors for $4000$ images for each of the $T=50$ diffusion timesteps. To measure the error solely for the analyzed step $t$, we start the inversion procedure from the exact latent from step (t1)(t-1) (cache from the sampling path). We split the latent pixels into plain and non-plain areas according to the masks calculated for clean images. More details on this setup can be found in Appendix J.
@@
In Fig. 4a, we present the visualization of calculated differences for each inversion step. There is a significant difference in the prediction errors for plain and non-plain areas, especially in the initial steps of the inversion process. Notably, for pixels associated with plain image areas, the error predominantly accumulates within the first 10% of the inversion steps. Additionally, in Fig. 4b, we present that this error discrepancy is strongly connected to a decrease in the diversity of diffusion models’ predictions. More precisely, we calculate a ratio of the predictions’ standard deviations
Figure 4: Discrepancy of the DDIM noise predictions for plain and non-plain image pixels. We show that, in the first inversion steps, the approximations are significantly (a) more erroneous and (b) less diverse for plain regions than for the rest of the image.
@@
between the sampling and inversion processes for the associated timesteps. We show that, for plain image regions, there is a significant decrease in the fraction of predictions’ variance preserved during the first inversion steps. Those observations can be related to recent works (Lee et al., 2023; Lin et al., 2024) analyzing why numerical solvers incur significant errors during the earliest diffusion steps (as $t\rightarrow 0$). Specifically, Lee et al. (2023) trace the error to the $1/t$ curvature blow-up of the reverse-time ODE trajectory, whereas Lin et al. (2024) attribute the predominance of early DDIM inversion approximations to a singularity arising from commonly used noise schedules. Our experiments extend those studies by showing that the error can be mainly attributed to the plain regions in the original images.
@@ Conclusion.
Early approximations during the DDIM Inversion procedure result in unequally distributed errors for pixels related to the plain and non-plain image areas, making it the origin of the structural patterns and correlations in the latents.
4 Consequences of the divergence and how to mitigate them?
After identifying the differences between the noises and latents, and highlighting the origin of this phenomenon, we finally pose the last question:
Research Question 4: What are the practical consequences of the divergence between noises and inverse DDIM latents, and how can they be mitigated?
Our findings in Section 3.3 indicate that the initial inversion steps predominantly contribute to the divergence between DDIM latent variables and Gaussian noise, in the form of insufficiently diverse approximations of diffusion model predictions. Therefore, as a simple fix to this issue, we propose to replace the first inversion steps with random noise, as in a forward diffusion process. The rationale behind this decision is twofold:
  • Substituting initial steps with Gaussian noise allows us to recover the noise variance exactly when the DDIM inversion fails to do so.
  • Recent studies (Deja et al., 2022; Liu et al., 2025; Li & Chen, 2024; Fesl et al., 2025) have shown that final steps of the backward diffusion do not contribute additional generative information, instead functioning as a data-agnostic denoising process. Therefore, exact inversion of those steps is less important from the perspective of accurate reconstruction.
The proposed inversion step is therefore defined as follows (see Appendix K for pseudocode):
$x_{t}=\begin{cases}\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,&\text{if }t\leq t^{\prime}\quad\text{(forward diffusion)}\\ \sqrt{\alpha_{t}}{x_{t-1}}+(\sqrt{1-\bar{\alpha}_{t}}-\sqrt{\alpha_{t}-\bar{\alpha}_{t}})\cdot\epsilon^{(t)}_{\theta}(x_{t-1},c),&\text{if }t>t^{\prime}\quad\text{(DDIM inversion)}\end{cases}\vskip-3.99994pt$ (5)
Before moving to practical applications, we first evaluate the effectiveness of this approach, with $N=10000$ images generated with $T=50$ steps using DiT and IF models. For such generations, we first noise them with a forward diffusion to the intermediate step $t^{\prime}$, followed by the DDIM inversion for $T-t^{\prime}$ steps, ending up with an approximation of the initial noise. In Table 4, we show that by replacing just $4\%$ of the inversion steps, we can completely remove correlations in the latents, up to the level of random Gaussian noise. To further evaluate this effect, in Table 5, we measure the size of the different parts of images (plain vs non-plain) after saving them with the PNG lossless compression. Compression is most effective in the parts related to plain images generated from the DDIM latents, which incline low diversity of their values. At the same time, replacing only 4% of inversion steps significantly reduces this issue. While our simple fix appears to effectively decorrelate the inversion latents, in the following sections, we showcase the consequences of the divergence between noise and latents in several practical use cases.
@@ 4 Consequences of the divergence and how to mitigate them?
forward ($\%T$) Corr. Absolute Error $\times 10^{-3}$ Corr. Absolute Error $\times 10^{-3}$
@@ 4.1 Interpolation quality
We start with the task of image interpolation, where the goal, for two given images, is to generate a sequence of semantically meaningful intermediary frames. Numerous methods (Dhariwal & Nichol, 2021; Song et al., 2021; Samuel et al., 2023; Zheng et al., 2024; Zhang et al., 2024b; Bodin et al., 2025) Diffusion Models with DDIM inversion technique, to calculate latents, interpolate them, and reconstruct the target image. Song et al. (2021) propose to use the spherical linear interpolation (SLERP, Shoemake (1985)), that, for two objects $x$ and $y$, with a coefficient $\lambda\in[0;1]$, is defined as $z^{(\lambda)}=\frac{\sin{(1-\lambda)\theta}}{\sin{\theta}}x+\frac{\sin{\lambda\theta}}{\sin{\theta}}y$, where $\theta=\arccos\left((x\cdot y)/(\lVert x\rVert\lVert y\rVert)\right)$.
Figure 5: The quality and diversity of images generated from interpolations of latents $\mathbf{\hat{x}_{T}}$} deteriorate along the path, as indicated by, accordingly, the FID peak and Recall decrease. In contrast, the quality of generations from noise $\mathbf{x_{T}}$} interpolations remains stable. Our simple fix, which is replacing 4% of the first inversion steps with forward diffusion, alleviates this issue.
In our experiment, we compare the quality of interpolations in the noise and latent spaces. To this end, we sample N=20N=20k noises, use DDIM with $T=50$ diffusion steps to generate images, and invert those images back into their latents. Next, we randomly assign pairs, which we interpolate with SLERP for $\lambda\in\{0,\frac{1}{6},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{5}{6},1\}$ and denoise. In Fig. 5, we show that, by calculating FID-10k, generations starting from interpolations between random noises (in orange) preserve consistent quality along the entire path – meaning that all interpolated images fall into the real images manifold. In contrast, this property collapses for the setting with DDIM latents as inputs (in purple). In the following rows, we show that worse interpolation results for latents stem from the decline in the variety of produced generations, as indicated by a lower recall (Kynkäänniemi et al., 2019), especially when getting closer to the middle of the interpolation path. Nevertheless, as presented in Fig. 5, using our fix for the first 4% of steps mitigates this issue, enabling higher-quality interpolations with better diversity of the intermediate points. In Section P.2, we present qualitatively that the proposed fix leads to more diverse interpolations, especially in the image background.
@@ 4.2 Diversity and quality of image edition
Apart from image interpolations, text-to-image diffusion models with DDIM inversion are often used for text-based edition, where a source image is first inverted and then reconstructed with a different target prompt (Hertz et al., 2022; Mokady et al., 2023; Garibi et al., 2024; Huberman-Spiegelglas et al., 2024). However, knowing that DDIM latents are less diverse in plain areas, we hypothesize that using them as a starting point might reduce the diversity and quality of the edited samples. To evaluate this, we use DiT (Peebles & Xie, 2023) and IF (StabilityAI, 2023) as conditional DMs with $T=50$ diffusion steps. For each model, we construct two sets of $1280$ randomly selected (1) source prompts $P_{S}$, used during the generation and inversion, and (2) target prompts $P_{T}$, used for edition. Using source prompts $P_{S}$, we generate images $I_{S}$ from Gaussian noise $\mathbf{x_{T}}$} and invert them back into the latents $\mathbf{\hat{x}_{T}}$}. Next, we regenerate images $\hat{I}_{T}$ from the latents, changing the conditioning to the target prompts $P_{T}$. We compare the edits with ground truth targets $I_{T}$ generated from the original noise $\mathbf{x_{T}}$} with $P_{T}$. In Fig. 6, we present the drawback of leveraging latents as starting points for the denoising, where the structures for $I_{S}$ images’ backgrounds limit editing performance in $\hat{I}_{T}$ target images.
Figure 6: Image editing by reversing the original image with a source prompt and reconstructing it with a target one. DDIM Inversion produces less diverse image changes when the manipulation is related to plain regions in source images, contrary to when using ground-truth Gaussian noise. Replacing the first inversion steps with forward diffusion leads to more editable latents.
@@
against $I_{S}$ LPIPS $\uparrow$ $0.59_{\pm 0.12}$ $0.56_{\pm 0.12}$ $8_{\pm 0.11} 0.33±0$$3_{\pm 0.11}$
@@
alignment CLIP-T ($P_{T}$) $\uparrow$ $0.681$ $0.662$ $0.353$ $0.614$
@@
Table 6: Diversity of editions (generations from noise $\mathbf{x_{T}}$} and latents $\mathbf{\hat{x}_{T}}$}, conditioned on target prompt) in relation to source images $I_{S}$ and their alignment with source and target prompts. The arrows ($\uparrow$$\downarrow$) indicate greater generation diversity and higher text alignment to the target prompt.
@@
In Table 6, we quantitatively measure this effect. First, we calculate the diversity of target generations ($I_{T}$$\hat{I}_{T}$) against source images ($I_{S}$). We use DreamSim distance (Fu et al., 2023), LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004), and cosine similarity of DINO features (Darcet et al., 2024) to measure the distance between two sets of generations. The experiment shows that edits resulting from latent encodings $\hat{I}_{T}$ are characterized by higher similarity (SSIM, DINO) and lower diversity (DreamSim, LPIPS) relative to starting images $I_{S}$ than the one resulting from noises $I_{T}$. At the same time, in the bottom rows of Table 6, we show that the correlations occurring in the latent encodings induce lower performance in text-alignment to target prompts $P_{T}$, which we measure by calculating cosine similarity between text embeddings and resulting image embeddings, both obtained with the CLIP (Radford et al., 2021) encoder.
@@
Figure 7: Replacing first DDIM inversion steps with forward diffusion improves editions’ pixel diversity (a) and prompt-alignment (b). For IF and DiT models, diversity of generations can be improved by leveraging the forward diffusion process in the first inversion steps, and denoising resulting latents with a different prompt. Additionally, we observe a boost in generations’ alignment to the target prompts, as indicated by the larger CLIP-T value for $P_{T}$ and the smaller one for $P_{S}$. We present details of the experiment in Appendix L.
@@
We further evaluate how replacing the first inversion steps with the forward diffusion affects the diversity and text alignment of generated outputs. As shown in Fig. 7a, swapping the first $4\%$ of DDIM inversion steps with forward diffusion improves the diversity of images generated from latents almost to the level of the samples from original noise. At the same time, as presented in Fig. 7b, replacing the first inversion steps leads to a significant decrease in generations’ alignment to the source prompt ($P_{S}$) and an increase in similarity to the target one ($P_{T}$). This can also be observed in visualizations (Fig. 6). Moreover, because of the random noise added as the initial inversion steps, as presented in Appendix M, our approach enables stochastic image editing, producing diverse manipulations of the same image.
4.3 Do existing inversion methods fix the correlations?
Finally, we investigate whether inversion methods introduced in prior work resolve the issue of selectively reduced latents diversity. To this end, we employ the Stable Diffusion XL (Podell et al., 2024) model, and using $2000$ generations from COCO-30K prompts, we measure resulting inversions’ normality, editability, and image reconstruction performance. Results in Table 7 indicate that methods based on predicted noise regularization, such as Pix2Pix-Zero (Parmar et al., 2023) and ReNoise (Garibi et al., 2024), while slightly improving the latents’ quality, do not offer significantly better editability, while being two times slower than DDIM. On the other hand, replacing DDIM approximation (Eq. 3) with reverse DPMSolver (Lu et al., 2022) leads to decorellated latents at the cost of high image reconstruction error. We show that our fix offers the best editability of latents with minimal reconstruction loss, all at the lowest computational cost. To be more precise, thanks to the fact that selected inversions steps replaced with the randomly sampled noise are the least important in terms of accurate reconstruction, we can observe no increase in the reconstruction error when replacing 2% of forward steps (up to the 2nd decimal point of MAE), while for 4% of steps the additional error is around 1% of pixel deviations - a value below the threshold usually employed by adversarial attacks as being not noticeable by a human eye (Madry et al., 2017).
Prior NFEs Normality Image Reconstruction CLIP Text Alignment Inv. time
Corr. $\downarrow$ KL $\times 10^{-2}\downarrow$ MAE $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ Source $\downarrow$ Target $\uparrow$ [s/image] $\downarrow$
Gaussian Noise $0.08_{\pm.01}$ 0.10 $31.88_{\pm 11.66}$ $73.34_{\pm 9.65}$
DDIM Inv. 50 $0.16_{\pm.02}$ 0.89 0.03 0.10±.05 27.58 $34.99_{\pm 11.36}$ $69.58_{\pm 10.17}$ $7.17_{\pm.01}$
Pix2Pix-Zero 50 $0.15_{\pm.02}$ 0.85 0.03 0.10±.05 27.35 $34.86_{\pm 11.39}$ $69.73_{\pm 10.12}$ $22.07_{\pm 1.84}$
ReNoise ($T=50,K=1$) 50 $0.14_{\pm.02}$ 0.73 0.04 0.09±.05 25.64 $34.89_{\pm 11.46}$ $69.87_{\pm 10.07}$ $19.86_{\pm.22}$
ReNoise ($T=25,K=2$) 50 $0.14_{\pm.02}$ 0.59 0.04 0.09±.04 24.81 $35.21_{\pm 11.61}$ $69.68_{\pm 10.09}$ $15.36_{\pm.51}$
ReNoise ($T=17,K=3$) 51 0.13±.02 0.47 0.06 $0.13_{\pm.05}$ 22.35 $35.79_{\pm 11.65}$ $69.04_{\pm 9.98}$ $14.31_{\pm.49}$
DPMSolver-1 ($T=50$) 50 0.09±.01 0.50 0.06 $0.30_{\pm.10}$ 22.55 $34.81_{\pm 11.40}$ $70.26_{\pm 10.42}$ $7.06_{\pm.00}$
DPMSolver-2 ($T=25$) 50 0.09±.01 0.26 0.06 $0.14_{\pm.07}$ 24.76 $34.69_{\pm 11.55}$ 71.24±9.91 $7.06_{\pm.00}$
Ours (forward 2%) 49 $0.14_{\pm.02}$ 0.71 0.03 0.10±.05 27.12 34.32±11.49 $70.21_{\pm 10.13}$ 7.00±.00
Ours (forward 4%) 48 0.09±.01 0.38 0.04 $0.14_{\pm.04}$ 25.68 $\mathbf{33.62}_{\pm 11.63}$ $\mathbf{72.17}_{\pm 9.94}$ 6.86±.01
Table 7: Evaluation of inversion methods across multiple metrics: latents normality, image reconstruction, prompt alignment, and speed. DDIM with the proposed fix offers a good trade-off between latent editability and image reconstruction, while increasing the inversion speed.